A bullet of mass 40kg is fired horizontally in to a wooden block with a speed of 10m/s the bullet is brought to rest in the block in 0.2s by constant resistance.
Calculate.
* Magnitude of the resistance.
*Distance moved by the bullet.
1
Expert's answer
2021-08-18T11:07:29-0400

Given:

"m=40\:\rm kg"

"v_i=10\:{\rm m/s}, \quad v_f=0\:\rm m/s"

"t=0.2\:\rm s"


(a) Magnitude of the resistance

"F=m\frac{\Delta v}{\Delta t}=40*\frac{10}{0.2}=4000\:\rm N"

(b) Distance moved by the bullet


"d=\frac{v_f+v_i}{2}t=\frac{10}{2}*0.2=1\:\rm m"